the conducting spheres of radius r1 and r2 are at the same potential then the ratio of their electric field just outside the sphere is?
exp wd the steps plz...

Grade:12

1 Answers

KJ S
11 years ago

Same potential implies
KQ1/R1=KQ2/R2

Ratio of electrical field at surface
(KQ1/R12)/(KQ2/R22)=R2/R1